OpenStudy (anonymous):

Which is the equation of the line that passes through (6,2) and is perpendicular to a line with slope -1/3?

OpenStudy (anonymous):

Which is the equation of the line that passes through (6,2) and is perpendicular to a line with slope -1/3? A) y-6=-1/3(x-2) B) y-2=1/3(x-6) C) y-2=3(x-6) D) y-6=-3(x-6) E) y-2=-3(x-6)